# The Flavor Composition of Supernova Neutrinos

### v2

@@ Abstract
Acknowledgments.— We thank Orlando Peres for useful discussions. AC thanks for the support received by the scholarships CAPES/PROEX No. 88887.511843/2020-00, CNPq No. 140316/2021-3, CAPES-PrInt No. 88887.717489/2022-00 and FAPERJ No. E-26/204.138/2022. The work of YP was supported by the São Paulo Research Foundation (FAPESP) Grant No. 2023/10734-3 and 2023/01467-1, and by the National Council for Scientific and Technological Development (CNPq) Grant No. 151168/2023-7. MMS thanks for the support provided by iTHEMS and ABBL from RIKEN, the N3AS Physics Frontier Center, and the NSF under cooperative agreement 2020275. This work would not have been possible without the collaborative environment provided by the Nordic Winter School on Multimessenger Astrophysics. We thank the event organizers, as well as the financial support by Markus Ahlers for AC to attend the event.
@@
    To describe the evolution of neutrinos in the region of standard matter effects, we need to determine the fraction of each matter eigenstate ($\nu_{1}^{m}$, $\nu_{2}^{m}$, and $\nu_{3}^{m}$) entering this region. To go from the flavor basis to the matter basis we rotate the subspace $\mu$-$\tau$ in $\rho_{\nu}^{F}$ (Eq. 5) by $\theta_{23}\approx 45^{o}$. For NO, the fractions are given by:
    $f_{\nu_{3}^{m}}^{NO}=\frac{1}{N}\sum_{k}\alpha_{k}^{2},$ (SM1)
@@
    $f_{\nu_{1}^{m}(\nu_{2}^{m})}^{NO}=\frac{1}{N}\sum_{k}^{N}\left(\frac{\beta_{k}^{2}+\gamma_{k}^{2}}{2}\pm\Re\left\{\beta_{k}\gamma_{k}e^{i\phi_{k}}\right\}\right).$ (SM2)
    Where $\alpha_{k}$, $\beta_{k}$, $\gamma_{k}$ and $\phi_{k}$ are given in Eq. 5 of the main text. Note that, due to the $\nu_{\mu}-\nu_{\tau}$ mixing in matter, the difference between the fractions of $\nu_{1}^{m}$ and $\nu_{2}^{m}$ is related to the off-diagonal elements of Eq. 5 as
    $\left|f_{\nu_{1}^{m}}^{NO}-f_{\nu_{2}^{m}}^{NO}\right|=\left|\frac{2}{N}\sum_{k}\Re\left\{\beta_{k}\gamma_{k}e^{i\phi_{k}}\right\}\right|$ (SM3)
@@
    In Eq. 5, we describe the density matrix of neutrinos emerging from self-induced conversions in the flavor basis. As a result, when the off-diagonal elements of Eq. 5 vanish, we can effectively describe the system as an ensemble of flavor eigenstates. Nevertheless, in the absence of a significant number of $\mu$ or $\tau$ leptons in the SN environment, there is nothing that distinguishes the flavor basis from any other basis in the $\mu$-$\tau$ sector. Therefore, we can rotate the $\mu$-$\tau$ sector to an arbitrary primed basis,
    $\rho^{\prime}_{\nu}=\frac{1}{N}\sum_{k=1}^{N}\left(\begin{array}[]{ccc}\alpha_{k}^{2}&0&0\\ 0&\beta_{k}^{\prime 2}&\beta_{k}^{\prime}\gamma_{k}^{\prime}e^{-i\phi^{\prime}_{k}}\\ 0&\beta_{k}^{\prime}\gamma_{k}^{\prime}e^{i\phi^{\prime}_{k}}&\gamma^{\prime 2}_{k}\end{array}\right)~{},$ (SM4)
    and assume that the off-diagonal terms in the primed basis vanish:
    $\frac{1}{N}\sum_{k=1}^{N}\beta^{\prime}_{k}\gamma^{\prime}_{k}e^{i\phi^{\prime}_{k}}\approx 0.$ (SM5)
@@
    $\rho_{\nu}^{F}=\frac{1}{N}\sum_{k=1}^{N}\left(\begin{array}[]{ccc}\alpha_{k}^{2}&0&0\\ 0&c_{\theta}^{2}\beta_{k}^{\prime 2}+s_{\theta}^{2}\gamma_{k}^{\prime 2}&\frac{1}{2}s_{2\theta}(\beta_{k}^{\prime 2}-\gamma_{k}^{\prime 2})\\ 0&\frac{1}{2}s_{2\theta}(\beta_{k}^{\prime 2}-\gamma_{k}^{\prime 2})&s_{\theta}^{2}\beta_{k}^{\prime 2}+c_{\theta}^{2}\gamma_{k}^{\prime 2}\end{array}\right)~{},$ (SM6)
    where $\theta$ parametrizes the rotation between the primed and flavor bases, with $s_{\theta}=\sin\theta$ and $c_{\theta}\equiv\cos\theta$. Here, we aim to demonstrate that even in such a situation, where destructive interference is known to occur in some other arbitrary primed basis, the off-diagonal terms of $\rho_{\nu}^{F}$ are suppressed. Since the primed basis is arbitrary, $2\theta$ can assume any value in the range $[-\pi/2,\pi/2]$ with equal probability ($s_{2\theta}$ in the range [-1,1]). Moreover, $X\equiv\frac{1}{N}\sum_{k=1}^{N}(\beta_{k}^{\prime 2}-\gamma_{k}^{\prime 2})$ can take any value within the range $[-1,1]$ (here, we assume $\sum\alpha_{k}^{2}=0$, concentrating all the probability in the $\mu$-$\tau$ sector). Thus, the question becomes: what is the probability that the absolute value of $\frac{1}{2}s_{2\theta}X$ exceeds a given value $Y$?
    To compute $P\left(\left|\frac{1}{2}s_{2\theta}X\right|>Y\right)$, we need to find $P\left(\frac{1}{2}s_{2\theta}X>Y\right)$ and $P\left(\frac{1}{2}s_{2\theta}X<-Y\right)$. In each instance, we need to separate the cases in which $X>0$ and $X<0$. Assuming $2\theta$ and $X$ are uniformly distributed, we obtain:
@@
    Now we can calculate the probability that the off-diagonal term in $\rho_{\nu}^{F}$ is below a certain benchmark value $Y$, assuming that destructive interference occurred in the primed basis. For instance, for $Y=0.2$, we find $P(\left|\frac{1}{2}s_{2\theta}X\right|<~{}0.2)=1-P(>0.2)\approx 0.66$, implying that 66% of the time (roughly $1$ standard deviation $\sigma$, in analogy to the normal distribution), the absolute value of the off-diagonal term is less than $0.2$ and peaks at zero (mean value).
    In light of Eq. SM3, and assuming $\alpha_{k}^{2}=0$ for simplicity, we find that
@@
    $=\left|s_{2\theta}X\right|\lesssim 0.4\hskip 14.22636pt.$ (SM10)
    These correspond to theoretical uncertainties in the calculated flavor ratios presented in the main text: we obtain $f_{\nu_{e}}^{NO}\lesssim 0.5+\sigma_{\nu_{e}}^{NO}$ (Eq. 10) with $\sigma_{\nu_{e}}^{NO}\in[-0.08,0.06]$, computed by varying $f_{\nu_{1}^{m}}^{NO}-f_{\nu_{2}^{m}}^{NO}$ within $[-0.4,0.4]$. Additionally, we obtain $f_{\nu_{e}}^{IO}\approx 1/3+\sigma_{\nu_{e}}^{IO}$ (Eq. 12) with $\sigma_{\nu_{e}}^{IO}\in[-0.1,0.1]$ when varying $f_{\nu_{1}^{m}}^{IO}-f_{\nu_{3}^{m}}^{IO}$ within $[-0.4,0.4]$.
    Note, however, that these uncertainties are conservative, as $\sum\alpha_{k}^{2}$ generally tends to be non-zero. Therefore, the theoretical uncertainties should have less impact on the final results than expressed here, since all the values in the $\mu$-$\tau$ sector of $\rho_{\nu}^{F}$ (including off-diagonal elements) should be smaller by a factor of $(1-\sum\alpha_{k}^{2})$. Additionally, the distribution of $X$ may be non-uniform and peaked towards zero, further reinforcing our results and potentially narrowing the ranges for $f_{\nu_{e}}$ even more (we will return to this point in item 4 below).
@@
    From the argument above, note that the suppression of the off-diagonal terms in the flavor basis -- and consequently the small value of $\left|f_{\nu_{1}^{m}}^{NO}-f_{\nu_{2}^{m}}^{NO}\right|$ (likely $\lesssim 0.4$) -- arises from the fact that, due to $\mu$-$\tau$ mixing in matter, there are many bases where decoherence can occur, and in each of these bases, there are certain parameter combinations for which the off-diagonal elements of Eq. SM6 are small. This should be contrasted with the situation where there is no $\mu-\tau$ mixing (for example, due to the presence of muons), and only the matter basis has physical significance:
    $\rho_{\nu}^{F}=\rho_{\nu}^{M}=\frac{1}{N}\sum_{k=1}^{N}\left(\begin{array}[]{ccc}\alpha_{k}^{2}&0&0\\ 0&\beta_{k}^{2}&0\\ 0&0&\gamma_{k}^{2}\end{array}\right)~{}.$ (SM11)
    Here, $M$ refers to the mass basis, which coincides with the flavor basis (F) in this context. In this case, $\left|f_{\nu_{1}^{m}}^{NO}-f_{\nu_{2}^{m}}^{NO}\right|=\left|X\right|=\frac{1}{N}\left|\sum_{k=1}^{N}(\beta_{k}^{2}-\gamma_{k}^{2})\right|$ is, in principle, uniformly distributed and has only a $40\%$ chance of being less than 0.4. Moreover, $\left|f_{\nu_{1}^{m}}^{NO}-f_{\nu_{2}^{m}}^{NO}\right|\gtrsim 0.6$ occurs $40\%$ of the time in the context of Eq. SM11, whereas in the presence of $\mu$-$\tau$ mixing, it would be only $17\%$ (for $\sum\alpha_{k}^{2}=0$).
@@
    If we include additional assumptions -- such as the improbability of extreme values of $\alpha_{k}^{2}$, or even $\beta_{k}^{\prime 2}$ and $\gamma_{k}^{\prime 2}$ in an arbitrary primed basis, which would imply that all neutrinos reaching the detector left the inner regions in the same state (an exception is made for the neutronization burst and early accretion phases, where $\sum\alpha_{k}^{2}\approx 1$ is known to occur and collective effects are negligible.) -- we can further narrow the range of $f_{\nu_{e}}^{NO}$, constraining it to the middle of the region described by Eq. 10, often resembling the value for IO in Eq. 12.
Antineutrinos.— For antineutrinos, the evolution Hamiltonian is similar to Eq. 1 (main text), with the modification $V_{e}\rightarrow-V_{e}$. Therefore, at $\rho_{m}\gg\rho_{m}^{H}$, $\bar{\nu}_{e}\approx\bar{\nu}_{1}^{m}$ for NO, and the possible fraction of electron antineutrinos on Earth is given by Eq. 10 (main text) with $\left|U_{e1}\right|^{2}$ in place of $\left|U_{e3}\right|^{2}$,
$f_{\bar{\nu}_{e}}^{NO}=\frac{1}{2}\left(1-\left|U_{e1}\right|^{2}\right)+\frac{a}{2}\left(3\left|U_{e1}\right|^{2}-1\right)~{}.$ (SM9)
@@
$\frac{1}{6}\lesssim f_{\bar{\nu}_{e}}^{NO}\lesssim\frac{2}{3}~{}.$ (SM10)
Eq. SM10 is applicable even in the presence of shock waves. On the other hand, for IO, $\bar{\nu}_{e}\approx\bar{\nu}_{3}^{m}$ above the $H$-resonance density, and $f_{\bar{\nu}_{e}}^{IO}$ is described by Eq. 10,
$f_{\bar{\nu}_{e}}^{IO}\lesssim 0.5~{}.$ (SM11)
If the $H$-resonance is non-adiabatic, transitions $\bar{\nu}_{3}^{m}\rightarrow\bar{\nu}_{1}^{m}$ modify the allowed regions for $f_{\bar{\nu}_{e}}^{IO}$, which will be given by Eq. SM9. Therefore,
$\frac{1}{6}\lesssim f_{\bar{\nu}_{e}}^{IO}(\text{non-ad $H$})\lesssim\frac{2}{3}~{}.$ (SM12)
In case antineutrinos emerge from the core as incoherent matter eigenstates, we have $f_{\bar{\nu}_{e}}\lesssim 0.7$ for both mass orderings. These results are depicted in Fig. SM1.
Figure SM1: Same as Fig. 1 (in the main text), but for antineutrinos.

### v3

@@ Abstract
Acknowledgments.— We thank Orlando Peres for useful discussions. A.C. thanks for the support received by the scholarships CAPES/PROEX No. 88887.511843/2020-00, CNPq No. 140316/2021-3, CAPES-PrInt No. 88887.717489/2022-00 and FAPERJ No. E-26/204.138/2022. The work of Y.P. was supported by the São Paulo Research Foundation (FAPESP) Grants No. 2023/10734-3 and No. 2023/01467-1, and by the National Council for Scientific and Technological Development (CNPq) Grant No. 151168/2023-7. M.M.S. is thankful for the support provided by iTHEMS and ABBL from RIKEN, the N3AS Physics Frontier Center, and the NSF under cooperative Agreement No. 2020275. This work would not have been possible without the collaborative environment provided by the Nordic Winter School on Multimessenger Astrophysics. We thank the event organizers, as well as the financial support by Markus Ahlers for A.C. to attend the event.
@@
    To describe the evolution of neutrinos in the region of standard matter effects, we need to determine the fraction of each matter eigenstate ($\nu_{1}^{m}$, $\nu_{2}^{m}$, and $\nu_{3}^{m}$) entering this region. To go from the flavor basis to the matter basis, we rotate the subspace $\mu$-$\tau$ in $\rho_{\nu}^{F}$ [Eq. (5)] by $\theta_{23}\approx 45^{o}$. For NO, the fractions are given by
    $f_{\nu_{3}^{m}}^{NO}=\frac{1}{N}\sum_{k}\alpha_{k}^{2}$ (A1)
@@
    $f_{\nu_{1}^{m}(\nu_{2}^{m})}^{NO}=\frac{1}{N}\sum_{k}^{N}\left(\frac{\beta_{k}^{2}+\gamma_{k}^{2}}{2}\pm\Re\left\{\beta_{k}\gamma_{k}e^{i\phi_{k}}\right\}\right),$ (A2)
    where $\alpha_{k}$, $\beta_{k}$, $\gamma_{k}$ and $\phi_{k}$ are given in Eq. (5) of the main text. Note that, due to the $\nu_{\mu}-\nu_{\tau}$ mixing in matter, the difference between the fractions of $\nu_{1}^{m}$ and $\nu_{2}^{m}$ is related to the off-diagonal elements of Eq. (5) as
    $\left|f_{\nu_{1}^{m}}^{NO}-f_{\nu_{2}^{m}}^{NO}\right|=\left|\frac{2}{N}\sum_{k}\Re\left\{\beta_{k}\gamma_{k}e^{i\phi_{k}}\right\}\right|$ (A3)
@@
    In Eq. (5), we describe the density matrix of neutrinos emerging from self-induced conversions in the flavor basis. As a result, when the off-diagonal elements of Eq. (5) vanish, we can effectively describe the system as an ensemble of flavor eigenstates. Nevertheless, in the absence of a significant number of $\mu$ or $\tau$ leptons in the SN environment, there is nothing that distinguishes the flavor basis $F$ from any other basis $B$ in the $\mu$-$\tau$ sector. Therefore, we can rotate the $\mu$-$\tau$ sector to another arbitrary basis,
    $\rho^{B}_{\nu}=\frac{1}{N}\sum_{k=1}^{N}\left(\begin{array}[]{ccc}\alpha_{k}^{2}&0&0\\ 0&\beta_{k}^{\prime 2}&\beta_{k}^{\prime}\gamma_{k}^{\prime}e^{-i\phi^{\prime}_{k}}\\ 0&\beta_{k}^{\prime}\gamma_{k}^{\prime}e^{i\phi^{\prime}_{k}}&\gamma^{\prime 2}_{k}\end{array}\right),$ (A4)
    and assume that the off-diagonal terms in this basis vanish:
    $\frac{1}{N}\sum_{k=1}^{N}\beta^{\prime}_{k}\gamma^{\prime}_{k}e^{i\phi^{\prime}_{k}}\approx 0.$ (A5)
@@
    $\rho_{\nu}^{F}=\frac{1}{N}\sum_{k=1}^{N}\left(\begin{array}[]{ccc}\alpha_{k}^{2}&0&0\\ 0&c_{\theta}^{2}\beta_{k}^{\prime 2}+s_{\theta}^{2}\gamma_{k}^{\prime 2}&\frac{1}{2}s_{2\theta}(\beta_{k}^{\prime 2}-\gamma_{k}^{\prime 2})\\ 0&\frac{1}{2}s_{2\theta}(\beta_{k}^{\prime 2}-\gamma_{k}^{\prime 2})&s_{\theta}^{2}\beta_{k}^{\prime 2}+c_{\theta}^{2}\gamma_{k}^{\prime 2}\end{array}\right),$ (A6)
    where $\theta$ parametrizes the rotation between the primed and flavor bases, with $s_{\theta}=\sin\theta$ and $c_{\theta}\equiv\cos\theta$. Here, we aim to demonstrate that even in such a situation, where destructive interference is known to occur in some other arbitrary basis $B$, the off-diagonal terms of $\rho_{\nu}^{F}$ are suppressed. Since the basis $B$ is arbitrary, we assume that $2\theta$ can assume any value in the range $[-\pi/2,\pi/2]$ ($s_{2\theta}$ in the range $[-1,1]$). Moreover, $X\equiv\frac{1}{N}\sum_{k=1}^{N}(\beta_{k}^{\prime 2}-\gamma_{k}^{\prime 2})$ can take any value within the range $[-1,1]$ (in order to maximize our uncertainty, we consider $\sum\alpha_{k}^{2}=0$, concentrating all of the probability in the $\mu$-$\tau$ sector).
    Thus, the question becomes: what is the probability that the absolute value of $\frac{1}{2}s_{2\theta}X$ exceeds a given value $Y$?
    To compute $P\left(\left|\frac{1}{2}s_{2\theta}X\right|>Y\right)$, we need to find $P\left(\frac{1}{2}s_{2\theta}X>Y\right)$ and $P\left(\frac{1}{2}s_{2\theta}X<-Y\right)$. In each instance, we need to separate the cases in which $X>0$ and $X<0$. Assuming $2\theta$ and $X$ are uniformly distributed, [^7] This assumption is made with the sole purpose of simplifying calculations, as our argument is merely a proof of principle; This can be straightforwardly generalized to different choices of probability distributions for $\theta$ and $X$. On the other hand, assuming that neutrinos follow completely different evolutions, states given by $X=\pm 1$ (i.e., pure states) should be much less likely than other combinations that lead to mixed states. This more realistic scenario, which deviates from the uniform distribution, would make our results even more robust but would also be more complicated to handle. More comments on this will be provided later. we obtain
@@
    Now we can calculate the probability that the off-diagonal term in $\rho_{\nu}^{F}$ is below a certain benchmark value $Y$, assuming that destructive interference occurred in the basis $B$. For instance, for $Y=0.2$, we find $P(\left|\frac{1}{2}s_{2\theta}X\right|<~{}0.2)=1-P(>0.2)\approx 0.66$, implying that 66% of the time (roughly $1$ standard deviation $\sigma$, in analogy to the normal distribution), the absolute value of the off-diagonal term is less than $0.2$ and peaks at zero (mean value).
    In light of Eq. (A3), and assuming $\alpha_{k}^{2}=0$ for simplicity, we find that
@@
    $=\left|s_{2\theta}X\right|\lesssim 0.4.$ (A10)
    These correspond to theoretical uncertainties in the calculated flavor ratios presented in the main text: we obtain $f_{\nu_{e}}^{NO}\lesssim 0.5+\sigma_{\nu_{e}}^{NO}$ [Eq. (10)] with $\sigma_{\nu_{e}}^{NO}\in[-0.08,0.06]$, computed by varying $f_{\nu_{1}^{m}}^{NO}-f_{\nu_{2}^{m}}^{NO}$ within $[-0.4,0.4]$. Additionally, we obtain $f_{\nu_{e}}^{IO}\approx 1/3+\sigma_{\nu_{e}}^{IO}$ [Eq. (12))] with $\sigma_{\nu_{e}}^{IO}\in[-0.1,0.1]$ when varying $f_{\nu_{1}^{m}}^{IO}-f_{\nu_{3}^{m}}^{IO}$ within $[-0.4,0.4]$.
    Note, however, that these uncertainties are conservative, as $\sum\alpha_{k}^{2}$ generally tends to be nonzero. Therefore, the theoretical uncertainties should have less impact on the final results than expressed here, since all the values in the $\mu$-$\tau$ sector of $\rho_{\nu}^{F}$ (including off-diagonal elements) should be smaller by a factor of $(1-\sum\alpha_{k}^{2})$. Additionally, the distribution of $X$ may be non-uniform and peaked toward zero, further reinforcing our results and potentially narrowing the ranges for $f_{\nu_{e}}$ even more (we will return to this point in item 4 below).
@@
    From the argument above, note that the suppression of the off-diagonal terms in the flavor basis--and consequently the small value of $\left|f_{\nu_{1}^{m}}^{NO}-f_{\nu_{2}^{m}}^{NO}\right|$ (likely $\lesssim 0.4$)--arises from the fact that, due to $\mu$-$\tau$ mixing in matter, there are many bases where decoherence can occur, and in each of these bases, there are certain parameter combinations for which the off-diagonal elements of Eq. (A6) are small. This should be contrasted with the situation where there is no $\mu-\tau$ mixing (for example, due to the presence of muons), and only the matter basis has physical significance:
    $\rho_{\nu}^{F}=\rho_{\nu}^{M}=\frac{1}{N}\sum_{k=1}^{N}\left(\begin{array}[]{ccc}\alpha_{k}^{2}&0&0\\ 0&\beta_{k}^{2}&0\\ 0&0&\gamma_{k}^{2}\end{array}\right).$ (A11)
    Here, $M$ refers to the mass basis, which coincides with the $F$ basis in this context. In this case, $\left|f_{\nu_{1}^{m}}^{NO}-f_{\nu_{2}^{m}}^{NO}\right|=\left|X\right|=\frac{1}{N}\left|\sum_{k=1}^{N}(\beta_{k}^{2}-\gamma_{k}^{2})\right|$ is, in principle, uniformly distributed and has only a $40\%$ chance of being less than 0.4. Moreover, $\left|f_{\nu_{1}^{m}}^{NO}-f_{\nu_{2}^{m}}^{NO}\right|\gtrsim 0.6$ occurs $40\%$ of the time in the context of Eq. (A11), whereas in the presence of $\mu$-$\tau$ mixing, it would be only $17\%$ (for $\sum\alpha_{k}^{2}=0$).
@@
    If we include additional assumptions--such as the improbability of extreme values of $\alpha_{k}^{2}$, or even $\beta_{k}^{\prime 2}$ and $\gamma_{k}^{\prime 2}$ in an arbitrary primed basis, which would imply that all neutrinos reaching the detector left the inner regions in the same state (an exception is made for the neutronization burst and early accretion phases, where $\sum\alpha_{k}^{2}\approx 1$ is known to occur and collective effects are negligible.)--we can further narrow the range of $f_{\nu_{e}}^{NO}$, constraining it to the middle of the region described by Eq. (10), often resembling the value for IO in Eq. (12).
Antineutrinos.— For antineutrinos, the evolution Hamiltonian is similar to Eq. (1) (main text), with the modification $V_{e}\rightarrow-V_{e}$. Therefore, at $\rho_{m}\gg\rho_{m}^{H}$, $\bar{\nu}_{e}\approx\bar{\nu}_{1}^{m}$ for NO, and the possible fraction of electron antineutrinos on Earth is given by Eq. 10 (main text) with $\left|U_{e1}\right|^{2}$ in place of $\left|U_{e3}\right|^{2}$,
$f_{\bar{\nu}_{e}}^{NO}=\frac{1}{2}\left(1-\left|U_{e1}\right|^{2}\right)+\frac{a}{2}\left(3\left|U_{e1}\right|^{2}-1\right).$ (A12)
@@
$\frac{1}{6}\lesssim f_{\bar{\nu}_{e}}^{NO}\lesssim\frac{2}{3}.$ (A13)
Equation (A13) is applicable even in the presence of shock waves. On the other hand, for IO, $\bar{\nu}_{e}\approx\bar{\nu}_{3}^{m}$ above the $H$-resonance density, and $f_{\bar{\nu}_{e}}^{IO}$ is described by Eq. (10),
$f_{\bar{\nu}_{e}}^{IO}\lesssim 0.5.$ (A14)
If the $H$-resonance is nonadiabatic, transitions $\bar{\nu}_{3}^{m}\rightarrow\bar{\nu}_{1}^{m}$ modify the allowed regions for $f_{\bar{\nu}_{e}}^{IO}$, which will be given by Eq. (A12). Therefore,
$\frac{1}{6}\lesssim f_{\bar{\nu}_{e}}^{IO}(\text{non-ad $H$})\lesssim\frac{2}{3}.$ (A15)
In case antineutrinos emerge from the core as incoherent matter eigenstates, we have $f_{\bar{\nu}_{e}}\lesssim 0.7$ for both mass orderings. These results are depicted in Fig. A1.
Figure A1: Same as Fig. 1 (in the main text), but for antineutrinos.